RC meeting and Run Plan, Friday Oct. 6, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)
Beam recovery and checkout to production after MCC beam study day
- 17:15 - Hall A Sweep completed and hall put into controlled access.
- 17;40 - Hall A in power permit. Requested TO to ramp up BB and SBS Magnets, rotate HH field to ON/ON, and perform an NMR.
- 18:02 - NMR: 41.6%
- 19:00 - SL phoned to say MCC can provide beam. Beginning Beam checkout procedure
- Took a while to get harp scan profile looking alright.
- 21:30 - 50k and 200k carbon hole plots look good. However GEM UVA layer 0 all tripped and not responding to trip reset procedure. Called GEM expert on call Saru to assist
- 21:30 - NMR: 44.7%
- 21:45 - GEM HV Issue resolved, request 5uA at no target position, with pol he3 xsof ysof values.
Continuous GEM HV Layer 0 trips while trying to observe BPM and IC at 5uA
- 22:30 - Saru reset GEM HV again and now it seems stable.
Begin slow ramp on Pol He3 target 5uA->45uA, observing BPM and IC rates
- 23:30 - Reached stable production running at 45uA! (beam a little more trippy that last night)
- 01;30 - BB UVA GEMs tripped again. SL unable to recover with reset trip button, python script or crate reset. Eventually Saru was able to get them back up by resetting multiple times. Shift crew instructed to do this for rest of shift untill we can investigate during the DAY.
- 08:20 - UVA GEMs tripped again on the new DAY SL. Again, repeated recovery efforts unsuccessful. Saru is investigatnig.
- 09:00 - Decided to turn Layer0 off for now and get beam back.
- 10:40 - Escorted access by GEM experts and Alex to diagnose and fix the issue, thanks everyone! In doing so restarted the power strip to PS,SH,Grinch,GEM. This caused issues with BBCAL HV which we had to diagnose. Thanks provakar + kate!
- 12:40 - Target and Detector issues all seem to be fixed. Back to beam permit, magnets up, HH coils rotated to ON/ON. >30min Down Beam checkout procedure.
- 13:20 - GEM HV issues again. Experts actively working on it. Meanwhile configuring DAQ for sbs gems.
- 15:01 - SBS GEM Config in daq configured, requesting beam and back to production
- No major accelerator problems today
- There will be a maintenance day during Day shift on Wednesday 10/11
- If you want to get into the hall, please communicate with the RC to get this organized!
Data collected :
- SWING: 0.133 C
- OWL: 0.684 C
- DAY: 0.265 C
- TOTAL: 1.082 (I USED A REAL CALCULATOR THIS TIME)
Hall A Subsystems Status
- Need to re-set the trip points for RTDs after alarm handler connection issues earlier.
- Target otherwise operating nominally
- Hodo HV readout bugged and displaying the VMon instead of VMon - VSet. Alarm Handler and HV GUI restarted many times today. HV values are fine though.
- BBCal 7.4 dead. Discussion on changing PMT base at next available time.
Hall access requests
If anyone wants to make an access, let the RC know!
Please check when your shifts are!
Please sign up for any open shifts!
Short Term Run plan
Hall B have asked when we next plan an IHWP change
- They and and IA charge asymmetry issue (related to LCW?) which is now resolved
- They need new values for IHWP in and out
- Therefore changing would allow them to get the out values.
- Standard production running at 45uA, with NMRs every 3 hours.
==> List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)